Can these questions be answered for my midterm review? Thank you!

1) Suppose a company purchased land and a building for $16,225,538 cash. The appraised value of the building was $12,355,025, and the land was appraised at $7,784,271.

What dollar amount of the purchase price will be allocated to the Building account?

*Note: for calculations, please do not round the percentage.

2) At the beginning of 2023, Apu purchases a new Squishee machine for the Kwik-E-Mart with a cost of $196,825. The new machine has an estimated life of 12 years or 97,585 units (squishee drinks). The estimated salvage value is $19,285. During 2023, 13,471 units (drinks) were made with the new machine and during 2024, 14,474 units (drinks) were made.

Calculate the accumulated depreciation on December 31, 2023, using units-of-production depreciation.

3) At the beginning of 2023, Apu purchases a new Squishee machine for the Kwik-E-Mart with a cost of $206,088. The new machine has an estimated life of 11 years or 102,383 units (squishee drinks). The estimated salvage value is $20,502. During 2023, 13,329 units (drinks) were made with the new machine and during 2024, 14,872 units (drinks) were made.

What is net book value at the end of 2025, using straight-line depreciation?
Answer #1

1) Total appraised value = 12355025+7784271 = 20139296

Dollar amount allocated to Building = 16225538/20139296*12355025 = 9954019

2) Accumulated depreciation on December 31,2023 = (196825-19285)/97585*13471 = 24508

3) Accumulated depreciation = (206088-20502)/11*3 = 50614

Net book value at the end of 2025 = 206088-50614 = 155474